Ahmed purchased car for AED 110,500 and then he rents it out for 5 years at AED 28,000 per year. If the annual maintenance cost for the car is AED 2,000 per year and the salvage value after 5 years is AED 70,000. Calculate the present worth value for this car investment at 15% interest rate.
                                                           

11,461

-12,023

20,876

-23,804
Answer #1

Option (1).

Annual net income = Rental income - Maintenance cost = 28,000 - 2,000 = 26,000

Present worth of car investment (AED) = - 110,500 + 26,000 x P/A(15%, 5) + 70,000 x P/F(15%, 5)

= - 110,500 + 26,000 x 3.3522** + 70,000 x 0.4972**

= - 110,500 + 87,157 + 34,804

= 11,461